OpenStudy (anonymous):

Solve the following equation by transforming it into a perfect square trinomial. 3x2 + 12x = 63 A. {–27, 23} B. {–29, 21} C. {–7, 3} D. {–9, 1}

OpenStudy (anonymous):

Factor the equation above: \[3(x ^{2}+4x-21)=0\] 3(x+7)(x-3)=0 solve for x

OpenStudy (anonymous):

x=-7 and x=3 C

OpenStudy (anonymous):

Thank you for explaination but I didnt get it. Thank you though! :)

hartnn (hartnn):

you know how to make a trinomial, a perfect square ? what steps have you tried yet ?

OpenStudy (anonymous):

this equation can be factored rather easily 3x^2 + 12x = 63 so in order to factor we need to set the equation equal to zero 3x^2+12x-63=0 all these numbers are divisible by 3 so we factor 3 out 3(x^2+4x-21)=0 so the factors of -21 that have the sum of 4 are 7 & -3 so 3(x+7)(x-3)=0 now we set the factors equal to zero x+7=0 x-3=0 x=-7 x=3 C Does this help a bit?

hartnn (hartnn):

solved any problem of completing the square ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so, first. bring it in the form of x^2+bc+c=0 could u?

hartnn (hartnn):

**x^2+bx+c=0